OpenStudy (anonymous):

Solve V = π r^2h for h is the answer h=v/πr^2?

OpenStudy (anonymous):

Given: \[V=\pi r^{2}h \] Isolate h:\[\frac{ V }{ h }=\frac{ \pi r^{2}h }{ h }\]\[\frac{ V }{ h }=\pi r^{2} \]\[h=\frac{ V }{ \pi r^{2} }\]